FOURIER TRANSFORM MICROWAVE SPECTRA OF n-BUTANOL AND ISOBUTANOL

\underline{TAIGO UZUYAMA}, YUGO TANAKA, and YOSHIYUKI KAWASHIMA, Department of Applied Chemistry, Faculty of Engineering, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN; EIZI HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

We have investigated two butanols, normal (1-butanol) and iso (2-methyl-1-propanol), by Fourier transform microwave spectroscopy, focusing attention mainly to internal motion and rotational isomerism. We anticipate that n-butanol exists in fourteen rotational isomers with the combination of the following conformations: trans (T) and gauche (G) with respect to the CH_{3}CH_{2}-CH_{2}CH_{2}OH bond, trans (t) and two gauche’s (g, g’) to the CH_{3}CH_{2}CH_{2}-CH_{2}OH bond, and two or three potential minima in the internal rotation of OH, and that isobutanol in five. We scanned the frequency region from 7 to 25 GHz using a heated nozzle, and detected six and three isomers for n-butanol and isobutanol, respectively: three with Ar as a buffer gas and three with Ne instead of Ar for n-butanol, whereas Ar was employed for isobutanol. The observed spectral lines were assigned by comparing the experimental moments of inertia and the observed spectral intensities with those calculated by ab initio MO at the MP2/6-311++G(d,p) level. We thus concluded that five conformers detected for n-butanol are T -form and one G-form and that three observed isomers of isobutanol are gauche with respect to the H(CH_{3})_{2}C-CH_{2}OH bond and one isomer lacking a-type spectra to trans to the CH_{2}-OH bond.